3 Marks Exercise 2: Given the following sens strand sequence to the Ubx bomeodomain: 5-cgaagacgcggccgacagacatacacccgctacca
Answer #1

Forward primer ACAGACATACACCCGCTAC

Reverse primer TTCTTCAGCTTCATTCGCC

Melting temperature formula : 4°C*(Number of G/C nucleotides) + 2°C*(Number of A/T nucleotides).

For forward primer :

4°C*(10) + 2°C*(9) = 40 + 18 = 58 °C

For reverse primer :

4°C*(9) + 2°C*(10) = 36 + 20 = 56 °C